OpenStudy (pinterestaddict15):

Will give medal. The polynomial x 3 + 5x 2 -57x -189 expresses the volume, in cubic inches, of a shipping box, and the width is (x+3) in. If the width of the box is 15 in., what are the other two dimensions? ( Hint: The height is greater than the depth.) A. height: 19 in. depth: 5 in. B. height: 21 in. depth: 5 in. C. height: 19 in. depth: 7 in. D. height: 21 in. depth: 7 in. I got this question wrong on a quiz, I thought it was C, but it was wrong. Does anyone know the correct answer?

OpenStudy (badhi):

Answer is B. We should be able to express the volume as following \[x^3+5x^2-57x-189 = (ax+b)(cx+d)(x+3)\] after finding values for a,b,c,d and substituting the value for x ( ie, its given that x+3 = 15) , you should get answer as B

OpenStudy (triciaal):